The table below pertains to Pieway, an economy in which the typical consumer’s basket consists of 10 bushels of peaches and 15 bushels of pecans.


  1. Table 24-1

    The table below pertains to Pieway, an economy in which the typical consumer’s basket consists of 10 bushels of peaches and 15 bushels of pecans.

    Year

    Price of
    Peaches

    Price of
    Pecans

    2005

    $11 per bushel

    $6 per bushel

    2006

    $9 per bushel

    $10 per bushel

       



    Refer to Table 24-1. The cost of the basket in 2006 was

     a.

    $210.

     b.

    $240.

     c.

    $245.

     d.

    $200.



  1. Table 24-1

    The table below pertains to Pieway, an economy in which the typical consumer’s basket consists of 10 bushels of peaches and 15 bushels of pecans.

    Year

    Price of
    Peaches

    Price of
    Pecans

    2005

    $11 per bushel

    $6 per bushel

    2006

    $9 per bushel

    $10 per bushel

       



    Refer to Table 24-1. If 2005 is the base year, then the CPI for 2005 was

     a.

    120.

     b.

    83.3.

     c.

    100.

     d.

    200.



  1. Table 24-1

    The table below pertains to Pieway, an economy in which the typical consumer’s basket consists of 10 bushels of peaches and 15 bushels of pecans.

    Year

    Price of
    Peaches

    Price of
    Pecans

    2005

    $11 per bushel

    $6 per bushel

    2006

    $9 per bushel

    $10 per bushel

       



    Refer to Table 24-1. If 2005 is the base year, then the CPI for 2006 was

     a.

    120.

     b.

    83.3.

     c.

    240.

     d.

    100.



  1. Table 24-1

    The table below pertains to Pieway, an economy in which the typical consumer’s basket consists of 10 bushels of peaches and 15 bushels of pecans.

    Year

    Price of
    Peaches

    Price of
    Pecans

    2005

    $11 per bushel

    $6 per bushel

    2006

    $9 per bushel

    $10 per bushel

       



    Refer to Table 24-1. If 2006 is the base year, then the CPI for 2005 was

     a.

    120.

     b.

    100.

     c.

    83.3.

     d.

    200.
Answer #1

Ans. Option b

Cost kf basket in 2006 = Total expenditure on peaches and pecans at 2006 prices = 10*9 + 15*10 = $240

Ans. Option c

Cost of basktet in 2005 = Total expenditure on peaches and pecans at 2005 prices = 10*11 + 15*6 = $200

CPI in 2005 with 2005 as base year = (Cost of basket in 2005/ cost of basket in 2005) *100 = 100

Ans. Option a

CPI in 2006 = (Cost of basket in 2006/Cost of basket in 2005)*100 = (240/200)*100 = 120

Ans. Option c

CPI of 2005 when 2006 is the base year = ( Cost of basket of goods in 2005/Cost if basket of goods in 2006)*100 = (200/240)*100 = 83.33